40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri May 16 10:54:42 2014, ericq, Update, LSC, X Arm ALS Noise coming in and out 
    Reply  Fri May 16 11:22:23 2014, Steve, Update, LSC, X Arm ALS Noise coming in and out 
Message ID: 9963     Entry time: Fri May 16 10:54:42 2014     Reply to this: 9964
Author: ericq 
Type: Update 
Category: LSC 
Subject: X Arm ALS Noise coming in and out 

Den and I spent some time with the interferometer last night with hopes of bringing in the AO path, but were stymied by the (re)occurrence of the anomalously high low frequency motion of the Xarm, as seen by fluctuations of TRX from .9 to .2 while "held" on resonance.

Jenne reported that they weren't seeing it earlier in the evening, and then it started again when I showed up. Holding the arms on IR, we could see a fair amount of excess low frequency noise in the BEATX_FINE_PHASE_OUT_HZ channel, as compared to BEATY, bringing its RMS to 5 times that of the Y arm. From the shape of the excess noise (broad slope from DC to tens of Hz), Rana suspected air currents and/or scattering effects being the culprit.

Den poked around a bit on the PSL table, which didn't really change much. He then went down to the X end table to inspect the table, and while he was there, I noticed the noise go down to being in line with the Yarm. I joined him at the end, and we found the beat phase noise in the frequency region of concern to be hugely sensitive to tapping on the enclosure, air current, etc. There is also a ton of green light everywhere, and multiple spots of green light around the green refl PD.

At that point however, the quiescent noise was acceptable (TRX fluctuations of <.2), so we went back to the control room to try to lock. Unfortunately, after a few attempts, the noise was back. At this point, we went home. The layout of the end table likely needs some attention to try and minimize our susceptibility to excess scatter effects. 

ELOG V3.1.3-